OpenStudy (anonymous):

Suppose g(x)=ax+b, where a and b are real constants. Find all possible pairs (a,b) such that we have g(g(x))=9x+28. Show work or explain.

myininaya (myininaya):

so g(x)=ax+b can you find g(g(x))?

OpenStudy (anonymous):

\[g(g(x)) =a ^{2}+ab+b\]

myininaya (myininaya):

do you mean \[a^2x+ab+b\]?

OpenStudy (anonymous):

yes

myininaya (myininaya):

but g(g(x))=9x+28 so \[9x+28=a^2x+(ab+b)\]

OpenStudy (anonymous):

\[a ^{2} x+ab+b=9x+28\]

myininaya (myininaya):

If this expression is equal the other expression then we must have the same amount of x's on both sides and the same amount of constant on both sides

myininaya (myininaya):